Use the driven equation x (t) from (C) and solve it using the following

numerical methods over the time interval 0 ≤ t ≤ 3 seconds at h=0.5.

I. Using the trapezium method

II. Using a Simpsons rule


1
Expert's answer
2021-05-23T23:35:01-0400

Let the driven equation be -

"x(t)=3t^2+2"


(i) Trapezoidal rule-


"=\int_0^3 (3t^2+2) dt" and h=0.5 assuming n=3


The values are-


"\int_0^3 (3t^2+2)dt=\dfrac{h}{2}(2+2.75+5+8.75)\\[9pt]

 =\dfrac{1.5}{2}\times 18.5
\\[9pt]
 =13.875"


Hence "\int_0^3 (3t^2+2) dt=13.875"


(ii) Using Simpson's rule-


    "=\int_0^3 (3t^2+2)dt\\[9pt]=\dfrac{3t^3}{3}+2t|_0^3\\[9pt]=(27-0)+(2(3-0)=33"